# countnullcountdimennullnullnulldimennullnullnulldimennullnullnulldimennullnullnulldimen

### v1

@@ 2.2 Distributionally robust revenue maximization
$X(\alpha)=\left\{\begin{array}[]{ll}x(\alpha)=\displaystyle\mu-\frac{d}{2\alpha}&\text{ w.p. }\alpha\\ y(\alpha)=\mu+\displaystyle\frac{d}{2(1-\alpha)}&\text{ w.p. }1-\alpha\end{array}\right$. (2.10)
for $\alpha\in[d/2\mu,1)$. When $X_{1},\dots,X_{m}$ are i.i.d. according to $\mathbb{P}_{\alpha}$, then $Y=\sum_{i=1}^{m}X_{i}\sim\mathbb{P}^{m}_{\alpha}$ with
3 Absolute revenue maximization
In this section we present the proofs of Theorem 1.1 and 1.3. One of the cornerstones to achieve our main results is Proposition 3.1 which might be of independent interest. Roughly speaking, it states that the sum of $m$ independent random variables with mean $\mu$ and mean absolute deviation (MAD) $d$ will be greater than or equal to $\approx m(\mu-d/2)$ with probability close to one as $m$ grows large. The proof of Proposition 3.1 is given in Appendix A.
Proposition 3.1 (One-sided concentration bound).
Let $\mu>0$ and $0<d<2\mu$ be given, and let $0<\epsilon<1-\frac{d}{2\mu}$. Then, for every $m\in\mathbb{N}$, it holds that
$\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\mathbb{P}^{m}\left(\sum_{i=1}^{m}X_{i}\geq(1-\epsilon)^{2}m\left(\mu-\frac{d}{2(1-\epsilon)}\right)\right)\geq 1-\frac{f(\mu,d,\epsilon)}{m},$ (3.1)
where $f(\mu,d,\epsilon)$ is a function independent of $m$.
3.1 Maximin analysis
In Lemma 3.2, we will show that the bundling mechanism yields a robust absolute revenue of $\mu-d/2$ for the maximin problem.
Lemma 3.2.
For $\mu>0$ and $0<d<2\mu$, it holds that
$\lim_{m\rightarrow\infty}\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}=\mu-\frac{d}{2},$ (3.2)
where the supremum is, roughly speaking, asymptotically attained by the bundling price
$p_{m}^{*}(\epsilon):=(1-\epsilon)^{2}m\left(\mu-\frac{d}{2(1-\epsilon)}\right),$ (3.3)
for $\epsilon>0$ sufficiently small.
Proof.
Recall from (2.2) that $\mathrm{BUND}(p_{m},\mathbb{P}^{m})=p_{m}\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m})$. Using Proposition 3.1, it follows that for every $0<\epsilon<1-d/2\mu$, we have
$\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{p_{m}\cdot\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m})}{m}$ $\geq\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{p_{m}^{*}(\epsilon)\cdot\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m}^{*}(\epsilon))}{m}$
$\geq\frac{p_{m}^{*}(\epsilon)}{m}\left(1-\frac{f(\mu,d,\epsilon)}{m}\right).$ (3.4)
The expression on the right approaches $p_{m}^{*}(\epsilon)/m$ as $m\rightarrow\infty$ for every fixed $0<\epsilon<1-d/2\mu$.
using that $\mathbb{P}^{m}(\sum_{i=1}^{m}X_{i}\geq p_{m})\leq 1$. Combining the lower and upper bounds in (3.4) and (3.5), respectively, it follows that for every fixed $\epsilon\in(0,1)$, we have
$(1-\epsilon)^{2}\left(\mu-\frac{d}{2(1-\epsilon)}\right)\leq\lim_{m\rightarrow\infty}\sup_{p_{m}}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\,\frac{\mathrm{BUND}(p_{m},\mathbb{P}^{m})}{m}\leq\mu-\frac{d}{2}.$
By letting $\epsilon\rightarrow 0$, we obtain the result in (3.2). ∎
3.2 Minimax analysis
In this section we show that for the minimax problem, and therefore also for the maximin problem, there does not exist a deterministic truthful mechanism that can do better than $\max\{\mu-d/2,d/2\}$.
Theorem 3.3.
Let $\mu>0$ and $0<d<2\mu$. There does not exist a deterministic truthful mechanism that can asymptotically achieve a (normalized with $m$) revenue greater than $\max\{\mu-d/2,d/2\}$ for the minimax problem. That is,
$\lim_{m\rightarrow\infty}\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\sup_{D\in\mathcal{D}_{m}}\frac{\,\mathrm{REV}(D,\mathbb{P}^{m})}{m}\leq\max\left\{\mu-\frac{d}{2},\frac{d}{2}\right\}.$ (3.6)
Proof.
We always have
$\inf_{\mathbb{P}^{m}\in\mathcal{P}(\mu,d)^{m}}\sup_{D\in\mathcal{D}_{m}}\,\mathrm{REV}(D,\mathbb{P}^{m})\leq\sup_{D\in\mathcal{D}_{m}}\,\mathrm{REV}(D,\mathbb{P}^{m}_{\alpha}),$
@@ 3.2 Minimax analysis
where we choose $\mathbb{P}^{m}_{\alpha}=\Pi_{i=1}^{m}\mathbb{P}_{\alpha}$ with $\mathbb{P}_{\alpha}$ the two-point distribution (2.10) with $\{x(\alpha(m)),y(\alpha(m))\}$ support and (parameterizing) probability $\alpha(m)=1-m^{-(m+1)}\cdot e^{-m}$. For a two-point distribution, every realized valuation for the goods is either $x$ or $y$. We will first argue that, based on the choice of $\alpha(m)$ and the individual rationality property, the only valuation vectors that matter for the revenue analysis, are those with at most one (realized) valuation being $y$. For a vector $v\in\{x,y\}^{m}$ we define $C_{y}(v)=|\{v_{i}:v_{i}=y\text{ for }i=1,\dots,m\}|$ to be the number of $y$’s in $v$. For any deterministic truthful mechanism $D$, we can then write the revenue as
$\frac{1}{m}\mathrm{REV}(D,\mathbb{P}^{m}_{\alpha})$ $=\frac{1}{m}\sum_{v:C_{y}(v)\leq 1}\pi(v)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$
$+\frac{1}{m}\sum_{k=2}^{m}\sum_{v:C_{y}(v)=k}\pi(v)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$
$\leq\frac{1}{m}\sum_{v:C_{y}(v)\leq 1}\pi(v)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$
$+\frac{1}{m}\sum_{k=2}^{m}\sum_{v:C_{y}(v)=k}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m})).$ (3.7)
@@
In the final inequality of (3.7) we use that, because of individual rationality, we always have $\pi(v)\leq\langle z(v),v\rangle\leq\sum_{i=1}^{m}v_{i}$. We first bound the second summation in (3.7). If $C_{y}(v)=k\geq 2$, we have, for $m$ large enough, that
$\frac{1}{m}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})$ $=(v_{1},\dots,v_{m}))=\frac{[(m-k)\cdot x(\alpha)+k\cdot y(\alpha)](\alpha)^{m-k}(1-\alpha)^{k}}{m}=$
$\leq C(\mu,d)\cdot m^{-(m+1)}\cdot e^{-m},$ (3.8)
where $C(\mu,d)$ is a constant only depending on $\mu$ and $d$ (independent of $k$ and $m$). The last inequality follows from the fact that the dominant term in the expression above is $(1-\alpha)^{k-1}$, using the definition of $x(\alpha)$ and $y(\alpha)$ in (2.10). Since $k\geq 2$, we then get the desired result as $(1-\alpha)^{k-1}\leq(1-\alpha)=m^{-(m+1)}\cdot e^{-m}$. This implies that, still for $k\geq 2$,
@@
$\frac{1}{m}\sum_{v:C_{y}(v)=k}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))\leq\binom{m}{k}C(\mu,d)\cdot m^{-(m+1)}\cdot e^{-m}\leq C(\mu,d)\cdot\frac{e^{-m}}{m}$
@@
using that $|v:C_{y}(v)=k|=\binom{m}{k}\leq m^{k}$. Finally, because $k\in\{2,\dots,m\}$ in the outer summation, i.e., there are m1m-1 terms, we obtain
$\frac{1}{m}\sum_{k=2}^{m}\sum_{v:C_{y}(v)=k}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))\leq(m-1)C(\mu,d)\cdot\frac{e^{-m}}{m}\leq C(\mu,d)\cdot e^{-m}.$ (3.9)
@@
$\frac{1}{m}\sum_{v:C_{y}(v)\leq 1}\pi(v)\mathbb{P}^{m}_{\alpha}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))=\frac{1}{m}\pi\left(v^{(0)}\right)\alpha^{m}+\frac{1}{m}\sum_{j=1}^{m}\pi\left(v^{(j)}\right)\alpha^{m-1}(1-\alpha),$
where $v^{(j)}$ is such that $v^{(j)}_{i}=\left\{\begin{array}[]{ll}y&\text{ for }i=j\\ x&\text{ for }i\neq j\end{array}\right$., and $v^{(0)}=(x,x,\dots,x)$ the vector only containing $x$’s. We write $\pi_{0}=\pi\left(v^{(0)}\right)$. We next argue that, without loss of generality, we may assume that $\pi\left(v^{(j)}\right)$ is the same for every $j=1,\dots,m$; we will denote their common value by $\pi_{1}$. [^4] In a nutshell, this follows from the fact that good identities are interchangeable in the definition of dominant strategy and individual rationality.
Claim 3.4.
Without loss of generality, we may assume that
  • $\pi\left(v^{(k)}\right)=\pi\left(v^{(\ell)}\right)=:\pi_{1}$ for all $k,\ell\in\{1,\dots,m\}$.
  • $z$ allocates the item with value $y$ to either all $v^{(j)}$ for $j=1,\dots,m$, or to none of them.

### v2

@@ 4 Proof of the Minimax Revenue Theorem
$\mathbb{P}_{N}=\left\{\begin{array}[]{ll}x_{0}=\mu-\frac{d}{2\alpha(m)}&\text{ w.p. }w_{0}=\alpha(m)=1-(Nm)^{-(m+1)}\cdot e^{-m}\\ x_{n}=\mu+\frac{d}{2N}\frac{m^{N}-1}{(m-1)m^{N-n}(1-\alpha(m))}&\text{ w.p. }w_{n}=\frac{(m-1)m^{N-n}}{m^{N}-1}(1-\alpha(m)),\textup{for }n=1,\dots,N\end{array}\right.$. (4.3)
To see that $\mathbb{P}_{N}\in\mathcal{P}(\mu,d)$, one can easily verify that $w_{n}\geq 0$, for all $n\in\{0,1,\dots,N\}$, and that the mean and MAD constraints are satisfied. Also,
$\sum_{n=1}^{N}w_{n}=(1-\alpha(m))\frac{(m-1)}{m^{N}-1}\sum_{n=1}^{N}m^{N-n}=(1-\alpha(m))\frac{(m-1)}{m^{N}-1}\sum_{k=0}^{N-1}m^{k}$
$=(1-\alpha(m))\frac{(m-1)}{m^{N}-1}\frac{m^{N}-1}{m-1}=1-\alpha(m),$
@@
so that $w_{0}+\sum_{n=1}^{N}w_{n}=1$. For this distribution, every realized valuation for the goods is either $x_{0}$, or $x_{1},\dots,x_{N}$. We will first argue that, based on the choice of $\alpha(m)$ and the individual rationality property, the only valuation vectors that matter for the revenue analysis, are those with at most one (realized) valuation being $x_{1},\dots,x_{N}$. For a vector $v\in\{x_{0},x_{1},\dots,x_{N}\}^{m}$ we define $C_{\alpha}(v)=|\{v_{i}:v_{i}>x_{0}\text{ for }i=1,\dots,m\}|$ to be the number of $x_{1},\dots,x_{N}$ in $v$. For any truthful mechanism $A$, we can then write the revenue as
$\frac{1}{m}\mathrm{REV}(A,\mathbb{P}^{m}_{N})$ $=\frac{1}{m}\sum_{v:C_{\alpha}(v)\leq 1}\pi(v)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$
$+\frac{1}{m}\sum_{k=2}^{m}\sum_{v:C_{\alpha}(v)=k}\pi(v)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$
$\leq\frac{1}{m}\sum_{v:C_{\alpha}(v)\leq 1}\pi(v)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$
$+\frac{1}{m}\sum_{k=2}^{m}\sum_{v:C_{\alpha}(v)=k}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m})).$ (4.4)
@@
In the final inequality of (4.4) we use that, because of individual rationality, we always have $\pi(v)\leq\langle z(v),v\rangle\leq\sum_{i=1}^{m}v_{i}$.
We first bound the term $\sum_{v:C_{\alpha}(v)=k}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))$ appearing in (4.4) for each $k\in\{2,\dots,m\}$. Fix $k\geq 2$. The set $\{x_{0},x_{1},\dots,x_{N}\}^{m}$ contains exactly $\binom{m}{k}N^{k}$ valuation vectors $v$ with $C_{\alpha}(v)=k$. Now fix any $v$ such that $C_{\alpha}(v)=k$. We can split $k=\sum_{n=1}^{N}k_{n}$ where $k_{n}$ denotes the number of coordinates $i\in\{1,\dots,m\}$ for which $v_{i}=x_{n}$. Define the index set $K=\{n\in\{1,\dots,N\}|k_{n}>0\}$. Then, for $m$ enough, we have that
$\frac{1}{m}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))=\frac{[(m-k)\cdot x_{0}+\sum_{n\in K}k_{n}\cdot x_{n}]w_{0}^{m-k}\prod_{n\in K}w_{n}^{k_{n}}}{m}$
$\leq[x_{0}+\sum_{n\in K}x_{n}]w_{0}^{m-k}\prod_{n\in K}w_{n}^{k_{n}}\leq x_{0}\prod_{n\in K}w_{n}^{k_{n}}+\sum_{n\in K}(x_{n}w_{n})w_{n}^{k_{n}-1}\prod_{l\in K\backslash\{n\}}w_{l}^{k_{l}}.$ (4.5)
We obtain this bound by using $(m-k)/m\leq 1$, $k_{n}/m\leq 1$, and $w_{0}=\alpha(m)\leq 1$. Next, for all $n\in K$, notice that $w_{n}^{k_{n}-1}\prod_{l\in K\backslash\{n\}}w_{l}^{k_{l}}\leq(1-\alpha(m))^{k_{n}-1}\prod_{l\in K\backslash\{n\}}(1-\alpha(m))^{k_{l}}\leq(1-\alpha(m))^{k-1}\leq 1-\alpha(m)$. Similarly, $\prod_{n\in K}w_{n}^{k_{n}}\leq 1-\alpha(m)$. Hence, (4.5) can be further bounded to
$\frac{1}{m}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))\leq[x_{0}+\sum_{n\in K}x_{n}w_{n}](1-\alpha(m))$
$\leq[x_{0}+\sum_{n=1}^{N}x_{n}w_{n}](1-\alpha(m))\leq 2\mu\cdot(Nm)^{-(m+1)}\cdot e^{-m}.$ (4.6)
The last inequality follows by observing that $x_{n}w_{n}\leq\mu+d/(2N)$. This implies that, still for $k\geq 2$,
@@
$\frac{1}{m}\sum_{v:C_{\alpha}(v)=k}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))\leq\binom{m}{k}N^{k}\cdot 2\mu\cdot(Nm)^{-(m+1)}\cdot e^{-m}\leq\frac{2\mu}{N}\cdot\frac{e^{-m}}{m}$
@@
using that $|v:C_{\alpha}(v)=k|=\binom{m}{k}N^{k}\leq m^{k}N^{k}\leq(Nm)^{m}$. Finally, because $k\in\{2,\dots,m\}$ in the outer summation, i.e., there are m1m-1 terms, we obtain
$\frac{1}{m}\sum_{k=2}^{m}\sum_{v:C_{\alpha}(v)=k}\left(\sum_{i=1}^{m}v_{i}\right)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))\leq(m-1)\frac{2\mu}{N}\cdot\frac{e^{-m}}{m}\leq\frac{2\mu}{N}\cdot e^{-m}.$ (4.7)
@@
$\frac{1}{m}\sum_{v:C_{\alpha}(v)\leq 1}\pi(v)\mathbb{P}^{m}_{N}((X_{1},\dots,X_{m})=(v_{1},\dots,v_{m}))=\frac{1}{m}\pi\left(v^{(0)}\right)w_{0}^{m}+\sum_{n=1}^{N}\frac{1}{m}\sum_{j=1}^{m}\pi\left(v^{(n,j)}\right)w_{0}^{m-1}w_{n},$ (4.8)
where $v^{(n,j)}$ is such that $v^{(n,j)}_{k}=\left\{\begin{array}[]{ll}x_{n}&\text{ for }k=j\\ x_{0}&\text{ for }k\neq j\end{array}\right$. and $v^{(0)}=(x_{0},x_{0},\dots,x_{0})$ the vector only containing $x_{0}$’s. We denote (averaged) payments as $\pi_{0}=\pi\left(v^{(0)}\right)$ and $\bar{\pi}_{n}=\frac{1}{m}\sum_{j=1}^{m}\pi(v^{(n,j)})$ for $n\in\{1,\dots,N\}$. We also define the average allocation probabilities as $\bar{s}_{0}=\frac{1}{m}\sum_{j=1}^{m}z_{j}(v^{(0)})$ and $\bar{s}_{n}=\frac{1}{m}\sum_{j=1}^{m}z_{j}(v^{(n,j)})$ for $n\in\{1,\dots,N\}$. Using these definitions, we can write (4.8) as
$\frac{1}{m}\pi_{0}w_{0}^{m}+\sum_{n=1}^{N}\bar{\pi}_{n}w_{0}^{m-1}w_{n}\leq\frac{1}{m}\pi_{0}+\sum_{n=1}^{N}\bar{\pi}_{n}w_{n}.$ (4.9)
By individual rationality, we have that
$\pi_{0}\leq\langle z(v^{(0)}),v^{(0)}\rangle=mx_{0}\bar{s}_{0}.$ (4.10)
Next, we use incentive compatibility regarding $v^{(n,j)}$ and $v^{(n-1,j)}$ for all $n\in\{1,\dots,N\}$. This results in
$\langle z(v^{(n,j)}),v^{(n,j)}\rangle-\pi(v^{(n,j)})\geq\langle z(v^{(n-1,j)}),v^{(n,j)}\rangle-\pi(v^{(n-1,j)}),\ \ j=1,\dots,m.$ (4.11)
We now expand both products. First, expanding the inner product on the left-hand side results in
$\langle z(v^{(n,j)}),v^{(n,j)}\rangle=\sum_{i\neq j}[x_{0}z_{i}(v^{(n,j)})]+x_{n}z_{j}(v^{(n,j)})=\sum_{i=1}^{m}[x_{0}z_{i}(v^{(n,j)})]+(x_{n}-x_{0})z_{j}(v^{(n,j)}).$ (4.12)
Then, expanding the inner product on the right-hand side results in
$\langle z(v^{(n-1,j)}),v^{(n,j)}\rangle=\sum_{i\neq j}[x_{0}z_{i}(v^{(n-1,j)})]+x_{n}z_{j}(v^{(n-1,j)})$
$=\sum_{i=1}^{m}[x_{0}z_{i}(v^{(n-1,j)})]+(x_{n}-x_{0})z_{j}(v^{(n-1,j)}).$ (4.13)
Hence, for all $j=1,\dots,m$, we obtain
$\pi(v^{(n,j)})\leq\pi(v^{(n-1,j)})+(x_{n}-x_{0})(z_{j}(v^{(n,j)})-z_{j}(v^{(n-1,j)}))+x_{0}\left[\sum_{i=1}^{m}z_{i}(v^{(n,j)})-\sum_{i=1}^{m}z_{i}(v^{(n-1,j)})\right]$
$\leq\pi(v^{(n-1,j)})+(x_{n}-x_{0})(z_{j}(v^{(n,j)})-z_{j}(v^{(n-1,j)}))+x_{0}m.$ (4.14)
Define $\Delta_{n}:=\bar{s}_{n}-\bar{s}_{n-1}$ for all $n\in\{1,\dots,N\}$. Summing (4) for each $j$ and dividing by $m$ gives
$\bar{\pi}_{n}\leq\bar{\pi}_{n-1}+(x_{n}-x_{0})(\bar{s}_{n}-\bar{s}_{n-1})+x_{0}m.$ (4.15)
We then obtain, recursively, by combining (4.15) and (4.10), the upper bound
$\bar{\pi}_{n}\leq nx_{0}m+mx_{0}\bar{s}_{0}+\sum_{k=1}^{n}(x_{k}-x_{0})\Delta_{k}\leq(N+1)m\left(\mu-\frac{d}{2}\right)+\sum_{k=1}^{n}(x_{k}-x_{0})\Delta_{k},$ (4.16)
and consequently,
$\sum_{n=1}^{N}\bar{\pi}_{n}w_{n}\leq N(N+1)m\left(\mu-\frac{d}{2}\right)(1-\alpha(m))+\sum_{n=1}^{N}w_{n}\sum_{k=1}^{n}(x_{k}-x_{0})\Delta_{k}\leq C(m)+\sum_{n=1}^{N}x_{n}w_{n}\Delta_{n}$ (4.17)
with $C(m)$ such that $\lim_{m\xrightarrow{}\infty}C(m)=0$. This follows because, for any $n\in\{1,\dots,N\}$, we have $\lim_{m\xrightarrow{}\infty}w_{n}x_{k}=\lim_{m\xrightarrow{}\infty}\left(w_{n}\mu+\frac{d}{2N}\cdot m^{-(n-k)}\right)=0$, for all $k<n$. We now focus on bounding $\sum_{n=1}^{N}x_{n}w_{n}\Delta_{n}$. We start by giving bounds on $\Delta_{n}$.
Lemma 4.1.
Let $n\in\{1,\dots,N\}$. The following holds:
$\Delta_{n}\geq 0.$